# Informed Correctors for Discrete Diffusion Models

### v4

@@ 5.2 Character Level Text Modeling
We next evaluate our method on the character-level modeling task using the text8 dataset (Mahoney, 2006). Following standard practice, we use the provided dataset splits and train on text chunks of length 256 for 1 million steps using a batch size of 512. As a simple measure of sample quality, we compute the error rate, defined as the fraction of total characters in words that are absent from the training set vocabulary.
We benchmark against MD4 (Shi et al., 2024), which employs a standard transformer architecture and has a comparable parameter count (fig.˜3). HollowDiff significantly outperforms the standard MD4 baseline in MD4 enhanced with the uninformed corrector and the ReMDM sampler, demonstrating the utility of our proposed sampling framework. We direct the readers to table˜3 for examples of generated text.
@@ 5.3 Class-conditional Image Modeling
We evaluate our method on tokenized ImageNet 256×256 using the VQ tokenizer from (Gu et al., 2022). All models are trained on the standard training split and evaluated without classifier-free guidance. Table 1 presents a comparison of generative models in the guidance-free setting. HollowDiff with the informed corrector achieves an FID of 6.26, using only 17 total steps (8 predictor, 8 corrector and a final denoising step). This outperforms several baselines with similar or larger model sizes and substantially more sampling steps. VQ-Diffusion (Gu et al., 2022) requires 100 steps to reach FID 11.89, while ReMDM (Wang et al., 2025) achieves FID 7.40 in 16 steps. We also note that the MD4 baseline achieves similarly strong results without any correctors, and the uninformed corrector and ReMDM sampler fail to improve its performance further. Generated samples from our method are shown in section˜D.3.4.
@@
To further demonstrate the value of informed correctors, we also test a hybrid setup where MD4 serves as the predictor and HollowDiff acts as the corrector. This combination improves FID to 5.78, exceeding the performance of both individual models, highlighting the versatility of the informed corrector even when paired with a stronger denoising predictor (see table˜1 for a full set of metrics). These findings confirm that our corrector provides complementary gains outside of architectural choice, and is broadly applicable to discrete diffusion models in general.
@@
Figure 4: FID vs. number of informed corrector updates $k$ under various sampling budgets.
We observe that for small NFE budgets, increasing $k$ significantly improves sample quality up to a point. For higher NFE budgets, performance is much less sensitive to $k$, and a small value (e.g., k=1k{=}1 or k=2k{=}2) already yields near-optimal FID. This aligns with the intuition that predictor steps dominate performance at high compute budgets. In all cases, too large $k$ (e.g., $k\geq 32$) leads to worse FID, suggesting interference among the simultaneous updates.
@@ A.1 Equivalence of objectives
For a complete picture, we include the derivation of the simplified objective in Shi et al. [2024], Sahoo et al. [2024], Ou et al. [2024] and show how it connects to the Campbell et al. [2022] objective expressed with CTMC transition rates. A similar treatment can be found in Shi et al. [2024, App. H.1]. The derivation highlights that while the two expressions are mathematical equivalent, they nonetheless exhibit important differences in when it comes to estimation cost.
@@ A.2 Simplifying the Score Function for Masked Diffusion
The relationship between the (concrete) score function and denoising distribution can be simplified in the case of masked diffusion. We first start with the relation between the denoising distribution and score function in  [Campbell et al., 2022]:
@@ Appendix B Informed Corrector
This translates to a straightforward procedure: find the non-mask dimensions with the lowest scores, and then sample from the model predictions $p^{\theta}_{0\mid t}(x^{d}\mid x^{\setminus d})$ of that dimension. For real-world datasets, we found that doing an argmax update in the end, where all dimensions are updated to the token that maximizes the conditional likelihood can be helpful for sample quality. We also follow Campbell et al. [2022] and apply the corrector only when $t<t_{c}$. We set entry time $t_{c}=0.9$ and number of correctors $C=1$ for the majority of our experiments. The algorithm is outlined in algorithm˜1 below.
@@
8:  if $t>t_{c}$ then
9:   Continue
10:  end if
11:  for $cstep=1$ to $C$ do
12:   Find most likely alternative tokens $x^{\prime d}=\arg\max_{x^{\prime d}\neq x^{d}}\log p^{\theta}_{0\mid t}(x^{\prime d}\mid x^{\setminus d})$
13:   Compute confidence score $c[d]=\log p^{\theta}_{0\mid t}(x^{d}\mid x^{\setminus d})-\log p^{\theta}_{0\mid t}(x^{\prime d}\mid x^{\setminus d})$, $\forall d\notin\mathcal{M}(x)$
14:   Sample Gumbel noise $g[d]\sim\text{Gumbel}(0,1)$
15:   Compute ranking criterion $r[d]=-c[d]+\tau g[d]$
16:   // Update the top-k dimensions
17:   for $cnt=1$ to $k$ do
18:    Get dimension of transition $d^{*}=\textsc{Sorted}(r)[D-cnt]$
19:    Update state $x^{d^{*}}\leftarrow\text{Categorical}(p^{\theta}_{0\mid t}(x^{d^{*}}\mid x^{\setminus{d^{*}}}))$
20:   end for
21:  end for
22:end for
23: // Deterministic update at the last step
24: Find most likely values for each mask dimension $x^{d}\leftarrow\mathop{\mathrm{argmax}}_{x^{d}_{0}}p^{\theta}_{0\mid t}(x^{d}_{0}\mid x)$ for all $d\in\mathcal{M}(x)$
25:Return $x$
@@ C.1 Hollow Transformer Architecture Details
We build upon the hollow transformer in [Sun et al., 2022] and design a class of hollow transformers that is more parameter efficient and expressive. First, we divide the computation into two streams, the content (causal) stream and the query (mixing) stream. The content stream consists of two transformers with causal self-attention in opposite directions, whereas the query stream attend to the content stream and outputs the final result. By design, non of the attention mechanisms allow a token at position $d$ to attend to a token that has access to information of $x^{d}$.
@@
where we insert a mixing layer for every $m$ forward/backward layers. Notably, the new mixing layer $M_{k}$ acts like a “query stream” while the forward and backward layers $F_{l}$ and $B_{l}$ act like “content streams” [Yang, 2019]. To prevent information leak and preserve the “hollowness,” we offset the inputs by one position when initializing the forward/backward layers:
@@ D.2 Text8
Figure 5: Loss curve on Text8 dataset Mahoney [2006]
@@ D.3.1 Training Details
Our model is adapted from that proposed by Sun et al. [2022] and consists of 24 standard transformer layers. We use an embedding dimension of 784 and hidden dimension of 2048. We add 3 mixture layers (1 after every 8 transformer layers) to the model to facilitate information exchange between the forward and backward streams. We train the model using AdamW Loshchilov and Hutter [2019] with an initial learning rate of 1e-4 and a batch size of 512. We use a linear learning rate warmup in the first 100 steps. We adopt a stepwise learning rate schedule, dropping the learning rate to 3.3e-5 at $2.11\times 10^{6}$ steps and to 1e-5 at $2.4\times 10^{6}$ steps. We stop the training at $2.5\times 10^{6}$ steps. The ImageNet experiments are done on v3-128 TPU pod machines on Google cloud and training takes 100 hours of wall clock time.
@@ D.3.3 Other Sampling Details
D.3.4 Generated Image Examples

### v5

@@ 5.2 Character Level Text Modeling
We next evaluate our method on the character-level modeling task using the text8 dataset [44]. Following standard practice, we use the provided dataset splits and train on text chunks of length 256 for 1 million steps using a batch size of 512. As a simple measure of sample quality, we compute the error rate, defined as the fraction of total characters in words that are absent from the training set vocabulary.
We benchmark against MD4 [12], which employs a standard transformer architecture and has a comparable parameter count (fig.˜3). HollowDiff significantly outperforms the standard MD4 baseline in MD4 enhanced with the uninformed corrector and the ReMDM sampler, demonstrating the utility of our proposed sampling framework. We direct the readers to table˜3 for examples of generated text.
@@ 5.3 Class-conditional Image Modeling
We evaluate our method on tokenized ImageNet 256×256 using the VQ tokenizer from [43]. All models are trained on the standard training split and evaluated without classifier-free guidance. Table 1 presents a comparison of generative models in the guidance-free setting. HollowDiff with the informed corrector achieves an FID of 6.26, using only 17 total steps (8 predictor, 8 corrector and a final denoising step). This outperforms several baselines with similar or larger model sizes and substantially more sampling steps. VQ-Diffusion [43] requires 100 steps to reach FID 11.89, while ReMDM [39] achieves FID 7.40 in 16 steps. We also note that the MD4 baseline achieves similarly strong results without any correctors, and the uninformed corrector and ReMDM sampler fail to improve its performance further. Generated samples from our method are shown in section˜D.3.5.
@@
To further demonstrate the value of informed correctors, we also test a hybrid setup where MD4 serves as the predictor and HollowDiff acts as the corrector. This combination improves FID to 5.78, exceeding the performance of both individual models (see table˜1 for a full set of metrics). Despite the hybrid setup using more total parameters (400M) than the individual HollowDiff or MD4 model (230M), it highlights the modularity and compatibility of our corrector: it can be applied to stronger pretrained predictors without need for joint training. This finding confirms that our corrector provides complementary gains outside of architectural choice, and is broadly applicable to discrete diffusion models in general.
@@
Figure 4: FID vs. number of informed corrector updates $k$ under various sampling budgets. Increasing $k$ initially improves efficiency but degrades sample quality for large $k$.
We observe that for small NFE budgets, increasing $k$ significantly improves sample quality up to a point. For higher NFE budgets, performance is much less sensitive to $k$, and a small value (e.g., k=1k{=}1 or k=2k{=}2) already yields near-optimal FID. This aligns with the intuition that predictor steps dominate performance at high compute budgets. When too many dimensions are updated simultaneously, their mutual dependencies are ignored, leading to update collisions where corrections can contradict or undo each other. This is observed empirically as in all cases, too large values of $k$ (e.g., $k\geq 32$) lead to worse FID.
@@ A.1 Equivalence of objectives
For a complete picture, we include the derivation of the simplified objective in Shi et al. [12], Sahoo et al. [16], Ou et al. [17] and show how it connects to the Campbell et al. [8] objective expressed with CTMC transition rates. A similar treatment can be found in Shi et al. [12, App. H.1]. The derivation highlights that while the two expressions are mathematical equivalent, they nonetheless exhibit important differences in when it comes to estimation cost.
@@ A.2 Simplifying the Score Function for Masked Diffusion
The relationship between the (concrete) score function and denoising distribution can be simplified in the case of masked diffusion. We first start with the relation between the denoising distribution and score function in  [8]:
@@ Appendix B Informed Corrector
This translates to a straightforward procedure: find the non-mask dimensions with the lowest scores, and then sample from the model predictions $p^{\theta}_{0\mid t}(x^{d}\mid x^{\setminus d})$ of that dimension. For real-world datasets, we found that doing an argmax update in the end, where all dimensions are updated to the token that maximizes the conditional likelihood can be helpful for sample quality. We set the number of correctors $C=1$ for the majority of our experiments. The algorithm is outlined in algorithm˜1 below.
@@
8:  for $cstep=1$ to $C$ do
9:   Find most likely alternative tokens $x^{\prime d}=\arg\max_{x^{\prime d}\neq x^{d}}\log p^{\theta}_{0\mid t}(x^{\prime d}\mid x^{\setminus d})$
10:   Compute confidence score $c[d]=\log p^{\theta}_{0\mid t}(x^{d}\mid x^{\setminus d})-\log p^{\theta}_{0\mid t}(x^{\prime d}\mid x^{\setminus d})$, $\forall d\notin\mathcal{M}(x)$
11:   Sample Gumbel noise $g[d]\sim\text{Gumbel}(0,1)$
12:   Compute ranking criterion $r[d]=-c[d]+\tau g[d]$
13:   // Update the top-k dimensions
14:   for $cnt=1$ to $k$ do
15:    Get dimension of transition $d^{*}=\textsc{Sorted}(r)[D-cnt]$
16:    Update state $x^{d^{*}}\leftarrow\text{Categorical}(p^{\theta}_{0\mid t}(x^{d^{*}}\mid x^{\setminus{d^{*}}}))$
17:   end for
18:  end for
19:end for
20: // Deterministic update at the last step
21: Find most likely values for each mask dimension $x^{d}\leftarrow\mathop{\mathrm{argmax}}_{x^{d}_{0}}p^{\theta}_{0\mid t}(x^{d}_{0}\mid x)$ for all $d\in\mathcal{M}(x)$
22:Return $x$
@@ C.1 Hollow Transformer Architecture Details
We build upon the hollow transformer in [33] and design a class of hollow transformers that is more parameter efficient and expressive. First, we divide the computation into two streams, the content (causal) stream and the query (mixing) stream. The content stream consists of two transformers with causal self-attention in opposite directions, whereas the query stream attend to the content stream and outputs the final result. By design, non of the attention mechanisms allow a token at position $d$ to attend to a token that has access to information of $x^{d}$.
@@
where we insert a mixing layer for every $m$ forward/backward layers. Notably, the new mixing layer $M_{k}$ acts like a “query stream” while the forward and backward layers $F_{l}$ and $B_{l}$ act like “content streams” [45]. To prevent information leak and preserve the “hollowness,” we offset the inputs by one position when initializing the forward/backward layers:
@@ D.2 Text8
Figure 5: Loss curve on Text8 dataset [44]
@@ D.3.1 Training Details
Our model is adapted from that proposed by Sun et al. [33] and consists of 24 standard transformer layers. We use an embedding dimension of 784 and hidden dimension of 2048. We add 3 mixture layers (1 after every 8 transformer layers) to the model to facilitate information exchange between the forward and backward streams. We train the model using AdamW [46] with an initial learning rate of 1e-4 and a batch size of 512. We use a linear learning rate warmup in the first 100 steps. We adopt a stepwise learning rate schedule, dropping the learning rate to 3.3e-5 at $2.11\times 10^{6}$ steps and to 1e-5 at $2.4\times 10^{6}$ steps. We stop the training at $2.5\times 10^{6}$ steps. The ImageNet experiments are done on v3-128 TPU pod machines on Google cloud and training takes 100 hours of wall clock time.
@@ D.3.3 Other Sampling Details
D.3.4 Ablation on confidence metrics.
We compared two confidence scoring schemes used in the informed corrector: the log-likelihood of the predicted token and the margin between the top-1 and top-2 logits. As shown below (table˜4), the margin-based confidence yields slightly better FID scores in both settings.
Model Confidence Type FID $\downarrow$
HollowDiff + informed corrector log-likelihood 6.45
HollowDiff + informed corrector margin 6.26
MD4 + informed corrector log-likelihood 5.85
MD4 + informed corrector margin 5.78
Table 4: Comparison of log-likelihood vs margin-based confidence scores.
D.3.5 Generated Image Examples